Find the complete integral of p3 +q^3 = 27z.


1
Expert's answer
2021-08-23T04:51:31-0400

Let "u= x+ay" where a is an arbitrary constant. Replacing p by "\frac{dz}{du}" and q by "a\frac{dz}{du}"

"(\frac{dz}{du})^3+(a\frac{dz}{du})^3= 27 z\\
(\frac{dz}{du})^3+a^3(\frac{dz}{du})^3= 27 z\\
(1+a^3)(\frac{dz}{du})=27z\\
\frac{1+a^3}{z}dz = 27u\\
Integrating \\
(1+a^3)Log \space z= 27 u +b\\
(1+a^3)Log \space z= 27 (x+ay) +b\\"